HUBBLE SPACE TELESCOPE DAILY REPORT #2097 PERIOD COVERED: 0000Z (UTC) 04/07/98 - 0000Z (UTC) 04/08/98 Daily Status Report as of 098/0000Z 1.0 ACTIVITIES SCHEDULED AND ACCOMPLISHED: 1.1 Completed STIS/CCD 7803 (Daily Darks to Update Acquisition Bad Pixel Table Part III) The Space Telescope Imaging Spectrograph (CCD) was used to make daily dark calibration updates. These daily darks are intended to identify hot pixels for update to the acquisition on-board FSW hot pixel table. Daily hot pixel lists can also be used for science data analysis. The observations were completed as planned with no anomalies. 1.2 Completed STIS/CCD 7926 (CCD Dark and Bias Monitor -- Continued) The Space Telescope Imaging Spectrograph (CCD) was used to make several dark and bias calibration observations. This is to measure CCD dark current and bias rates for the two supported amplifier settings {gain = 1 and gain = 4} in order to gather data for making superdarks and superbiases. These observations will also track the growth of hot pixels on the CCD in general and within the various target acquisition apertures specifically. The observations were executed as scheduled with no anomalies were. 1.3 Completed WF/PC-2 7712 (WF/PC-2 Cycle 7 Supplemental Darks pt2/3) The WF/PC-2 was used to make a series of Cycle 7 supplemental calibration darks. This dark calibration program obtains three dark frames every day to provide data for monitoring and characterizing the evolution of hot pixels. The observations were executed as scheduled, and no anomalies were noted. 1.4 Completed Three Sets of NIC/1/2/3 7850 (Cosmological Parameters Omega and Lambda from High-Z Type Ia Supernovae) The Near Infrared Camera (Cameras One, Two and Three) was used to make observations of high Z type Ia Super Novae 14H0.9Z2, 11H0.9Z1 and 14H0.9Z1 in order to measure the Omega and Lambda parameters. Type Ia supernovae can be used as calibrated standard candles to measure large distances. About one-half hour before the end of this six-and-a-quarter-hour first iteration of this proposal, the NICMOS suspend described in 2.1 occurred. The second iteration was not executed due to the suspend. Subsequently, NICMOS was recovered, and the final iteration was performed nominally. The observations were completed as planned, and no anomalies were reported. 1.5 Completed Five Sets of STIS/CCD 7908 (POMS Test Proposal: STIS Non-Scripted Parallel Proposal) The Space Telescope Imaging Spectrograph (CCD) was used to make calibration observations as part of the POMS Test Proposal. This was a STIS non-scripted parallel proposal. The observations were completed as planned, and no problems were reported. 1.6 Completed Two Sets of WF/PC-2 7909 (POMS Test Proposal: WF II Parallel Archive Proposal) The WF/PC-2 was used to make parallel observations as part of the POMS test proposal. This test proposal is designed to simulate scientific plans. The observations were completed as planned, and no problems were reported. 1.7 Completed WF/PC-2 7385 (Low-Mass Halo Binaries) The WF/PC-2 was used to make observations of the low mass, halo, binary, {Population II}, M-type, sub-dwarf, star LHS320. This star is too faint to be resolved through ground-based speckle interferometry, thus HST represents the only means of probing separations. The observations were completed as planned, and no anomalies were reported. 1.8 Completed NIC/1/2 7218 (Nearby Starburst Galaxies) The Near Infrared Camera (Cameras One and Two) was used to make observations of the nearby starbursting galaxy M-82. This proposal encompasses a study of both the stars and interstellar material in starburst nuclei which will test ideas about the triggering of starbursts and whether starbursts propagate outwards in a nucleus. The observations were completed as planned, and no anomalies were reported. 2.0 FLIGHT OPERATIONS SUMMARY: 2.1 Guide Star Acquisitions: Scheduled Acquisitions : 5 Successful : 5 Scheduled Reacquisitions : 10 Successful : 10 2.2 FHST Updates: Scheduled : 10 Successful : 10 2.3 Operations Notes: As documented in HSTAR 6429, the NICMOS suspended at 097/110304Z due to an "Intel Exception". The first two iterations of the proposal described in 1.4 were affected by this event. Using ROP NS-9, a NICMOS memory dump was performed at 097/1136Z. The dump was analyzed by Code 582. The NICMOS was successfully recovered at 097/165330Z per operations request. At approximately the same time, using ROP NS-5, the NSSC-1 error counter was cleared. Normal NICMOS science resumed at 097/1723Z. The SSR EDAC error counter was cleared on two occasions by ROP SR-1. For the current SMS, PASS software erroneously maneuvers the HGA during an on-going TDRSS event. This occurs two times on day 098. HSTAR 6430 was written. 3.0 SIGNIFICANT FORTHCOMING EVENTS: Continuation of normal science observations and calibrations. /DMH